OpenStudy (anonymous):

what is the product of 2(sqrt)9d (times) 3d(sqrt)5

OpenStudy (mertsj):

\[2\sqrt{9d}(3d \sqrt{5})\] \[6\sqrt{d}(3d \sqrt{5})=18d \sqrt{5d}\]